Calculate the standard free energy change for the reaction
H2 + I2 -------> 2HI at 127°c
Given that Kp= 41 atmosphere. ( R= 8.314 JK^-1 mol-1)
1
Expert's answer
2020-02-06T06:28:11-0500

Task:

Calculate the standard free energy change for the reaction

H2 + I2 ---> 2HI at 127°C

Given that Kp = 41 atmosphere. ( R = 8.314 J K-1 mol-1)


Solution:

ΔGo = standard free energy change.

The free energy change for a process taking place with reactants and products present under nonstandard conditions (ΔG) is related to the standard free energy change (ΔG°) according to this equation:

ΔG = ΔGo + RTln(Q)

Q is the reaction quotient.

R = gas constant = 8.314 J K-1 mol-1

T = 127°C = 400 K.

For a system at equilibrium, Q = Kp and ΔG = 0, and the previous equation may be written as:

0 = ΔGo + RTln(Kp) (at equilibrium)

ΔGo =−RTln(Kp).

Hence,

ΔGo = - 8.314 * 400 * ln(41) = -12349.85 J/mol = -12.35 kJ/mol.


Answer: ΔGo = -12.35 kJ/mol.